Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
carbognani - 4:04 Wednesday 22 March 2023 (59365) Print this report
Comment to Operator Report - Afternoon shift (59363)

Measurements have ended at iteration 215 (instead of completing all 225) with the same message from the post-processing library regarding finding centroid. Collection of "too dark" image is once again to be suspected. Experts should confirm.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×